OpenStudy (anonymous):

Differentiate y = 4loge(6 - 3x)

OpenStudy (lgbasallote):

\[\large y = 4\log_e (6-3x)\] is that it

OpenStudy (lgbasallote):

i mean is that the question

OpenStudy (anonymous):

yeah it is

OpenStudy (ash2326):

Use chain rule \[\frac{dy}{dx}=\frac{d}{dx} (4 \ln (6-3x))\] \[=4\frac{1}{(6-3x)}\frac{d}{dx} (6-3x)\] I think you can take it from here, can't you?

OpenStudy (anonymous):

I get to